User Tools

Site Tools


uc:uc20250902_september_2_2025

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

uc:uc20250902_september_2_2025 [2025/08/26 22:01] (current)
nchanove created
Line 1: Line 1:
 +==== APO 3.5-m Users Committee Meeting, 09/02/2025 ====
 +
 +----
 +
 +UPDATE EVERYTHING BELOW THIS:
 +
 +Attending: Nancy Chanover (NMSU), Bill Ketzeback (APO), Sarah Tuttle (UW), Chip Kobulnicky (UWy), Ben Williams (UW), Michael Hayden (OU), Misty Bentz (GSU), Eric Bellm (UW), Moire Prescott (NMSU)
 +
 +
 +=== User feedback and comments from institutional representatives ===
 +
 +  * NAPG - no report
 +  * UVa - nothing to report (by email)
 +  * Oklahoma - nothing to report
 +  * Washington - nothing to report
 +  * Seattle - nothing to report (by email)
 +  * NMSU - There is still an interest in trying to understand/mitigate the abnormally high background seen when taking long (20-30 min) exposures with KOSMOS. It manifests as a snow pattern in the background pixels. The engineering staff is still working to address this issue but if other users have seen this in their data it would be really helpful to know.
 +  * JHU - nothing to report (by email)
 +  * Colorado - nothing to report (by email)
 +  * Georgia State - nothing to report
 +  * Wyoming - nothing to report
 +  * BYU - nothing to report (by email)
 +
 +----
 +
 +=== Telescope and Instruments Report === 
 +
 +The detailed site report is included below, followed by additional information discussed during today's meeting.
 +
 +// 3.5-m Telescope and Instruments Highlights, 7/02/25 – 08/05/25 //
 +
 +1) Overview
 +
 +Summer shutdown activities occupied the first half of the month. Surprisingly in the second half of July we were able to get several nights of on-sky science despite the monsoon weather.
 +
 +Reminder: We have position openings for 3.5m Chief Telescope Technologist (Engineer) and Mechanical Engineer, Lead. These positions close on August 15 so please spread the word. 
 +
 +Mechanical Engineer: [[https://careers.nmsu.edu/jobs/mechanical-engineer-ld-remote-locations-new-mexico-united-states]]
 +
 +Chief Telescope Technologist: [[https://careers.nmsu.edu/jobs/chief-telescope-technologist-remote-locations-new-mexico-united-states]]
 +
 +We will be hiring only one person but have two postings to reflect different academic degrees and experience levels; interested candidates should apply for the one that best matches their skill set and preparation.
 +
 +2) Operations
 +
 +3.5m Telescope: Telescope is working as expected. Seasonal motion errors have been more frequent. The tertiary rotation mechanism has experienced some issues after a software remapping of port positions. Further position refinements may be required. 
 +
 +0.5m Telescope: Telescope is working as expected. Users have experienced several camera problems with both FlareCam and dcamera-spare related to frosting/icing or inability to cool properly. Cary serviced dcamera-spare by replacing both cooling fans and recharging the CCD vessel with gas. The off-axis guider has not been very reliable, requiring power cycles. 
 +
 +KOSMOS: System is cooled and stable. Increased dark current for long exposures has been confirmed. We attempted a vacuum servicing of the cryostat but without much luck. We anticipate having to open the vacuum vessel in the clean room sometime during Quarter 4 2025 to investigate possible thermal shorts or opens. kcamera-ICC was backed up over shutdown. 
 +
 +ARCTIC: The diffuser rotation mechanism is still unreliable even after a full servicing. Troubleshooting is continuing. The mechanism that moves the diffuser in and out of the optical path is still functional in the meantime. The rest of the instrument is cooled. We replaced the ion pump cable and moved the controller back to the intermediate level.
 +
 +Agile: The camera is non-operational; the thermoelectric cooler failed again. The camera is warm and we do not plan further repair work with this camera. We are planning on decommissioning the camera. The Agile instrument rotator is still not performing nominally and we are troubleshooting it further to prepare for SoonerCam.
 +
 +ARCES: The CCD reservoir for the cooling system is currently warm and not usable for science. We are keeping the shield reservoir cold. During shutdown we did a vacuum servicing but shortly after cooling the inter-order light showed elevated contamination. We have taken it out of service to address this problem. Work continues on a replacement and modern ICC, and so far testing is going well. A commissioning report for the new ICC is being worked on. 
 +
 +DIS: System is warm and out of service. Decommissioning has begun.
 +
 +NICFPS: System is cooled and usable. The ICC was backed up fully during shutdown.
 +
 +TripleSpec: System is cooled and usable. The ICCs (tspec and tcamera) were backed up during shutdown. Tcamera-ICC experienced another failure of one of its RAID hard drives and it needed to be rebuilt. 
 +
 +APOLLO: A team from NASA returned to site this past month to address issues with optical alignment and laser power issues. The instrument is again available for laser ranging but additional work will need to be done upgrading the power measurement equipment and working on some of the thermal metrology.
 +
 +All-Sky camera: 3.5m engineering staff has been working with the ALCOR manufacturer to remedy camera and software problems. Troubleshooting continues.
 +
 +----
 +
 +=== Additional site, telescope and instruments discussion === 
 +
 +xxx add stuff here xxx
 +
 +----
 +
 +=== 2025 Q3 3.5m scheduling ===
 +
 +The 3.5m schedule for Q3 has been in flux given the recent echelle engineering. Users who lost observing time due to the echelle servicing or weather are especially encouraged to request OPEN/DD time in Q3. To request this time please follow the standard procedure by emailing your request to Ben Williams, Russet McMillan, Amanda Townsend, Nancy Chanover, and your institutional scheduler.
 +
 +----
 +
 +=== 2025 Q4 3.5m scheduling ===
 +
 +All allocations for Q4 have been sent to the institutional schedulers; proposals are due on August 25. Note that the allocation emails stated that our ability to offer a virtual training in Q4 is uncertain due to staffing limitations; we indicated that we MAY offer it in Q4.
 +
 +----
 +
 +=== 2025 Q3 0.5 scheduling ===
 +
 +Next week is the only open week remaining in the ARCSAT Q3 schedule.
 +
 +----
 +
 +=== Status of new instrument initiatives === 
 +
 +  * SoonerCam: we received a modest number of survey responses and will proceed using the information we have. The camera was purchased and a meeting is scheduled in two weeks to discuss the next steps.
 +  * KOSMOS grisms: we are still awaiting a decision on the proposal we submitted for funds to support the purchase of three new grisms for KOSMOS.
 +  * ASPEN: we are still awaiting a decision on the proposal we submitted for funds to support the development of the new echelle spectrograph. Work on the design of the new optical echelle spectrograph continues with the goal of completing the Preliminary Design Review in early 2026.
 +  * Ocotillo: a number of purchase orders were executed and we expect the first batch of hardware to come in within the next month or so.
 +
 +----
 +
 +=== ACTION ITEMS === 
 +
 +Open action items from previous meetings:
 +  * UC reps: ask your users to complete the survey from FAWG (per Nancy’s email sent on 7/23/25). Status: CLOSED.
 +
 +
 +New action items from this meeting:
 +  * None.
 +
 +----
 +
 +=== All Other Business ===
 +
 +None.
 +
 +----
 +
 +=== Next meeting ===
 +
 +The next meeting will be on October 7, 2025 at 10:30 MDT.
 +
 +----
  
uc/uc20250902_september_2_2025.txt · Last modified: 2025/08/26 22:01 by nchanove